OpenStudy (anonymous):

@whpalmer4 1. express as the sum of difference of logarithms log6 11y 2. D(x)=10 log(10^16 x) if the intensity c of a sound changes by a factor of 1000 by how much does decibel level change? increase or decrease by how many decibels?

OpenStudy (mathmale):

Could you possibly share your "1. express as the sum of difference of logarithms log6 11y" by drawing it, using the Draw utility (below)? Do you mean "the logarithm to the base 6 of 11y" Once I'm clear on what you have to work with, I can give you better feedback.

OpenStudy (whpalmer4):

I think that's \[\log_6 11y\]Property of logs in question here is \[\log a + \log b = \log a*b\]

OpenStudy (whpalmer4):

Is the second formula \[D(x) = 10\log(10^{16} x)\]? That seems a bit off to me...